OpenStudy (anonymous):

HELP FOR A MEDAL AND BEST ANSWER REALLY STUCK ON THIS :// you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (agent0smith):

To start off: The function f(x) is a difference of squares means something like x^2 - 4 for a sum of squares, change that minus sign to a plus sign

OpenStudy (anonymous):

ok c:

OpenStudy (agent0smith):

Still need more help?

OpenStudy (anonymous):

i dont really get it tbh

OpenStudy (agent0smith):

Don't really get what? What i wrote, or the assignment?

OpenStudy (agent0smith):

@omgkelley

OpenStudy (agent0smith):

The function f(x) is a difference of squares. This means a function of the form \[\Large f(x) = x^2 - a^2\]they're both squares, and the difference means subtraction. For example \\[\Large f(x) = x^2 -9\]since 9 is 3 squared.

OpenStudy (anonymous):

oh okay c: & the assignment

OpenStudy (agent0smith):

That's why we're doing it step by step... now do you know how to factor this: \[\Large f(x) = x^2 -9\]